OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (uri):

Solve the parentheses first...What do you get? @BSANDSTORM

OpenStudy (uri):

−3*(x + 3) = −3*(x + 1)

OpenStudy (anonymous):

-3x+-9=-3x+-3-5

OpenStudy (uri):

+ -=-

OpenStudy (uri):

-3x-9=-3x-3-5 right?

OpenStudy (anonymous):

what's next

OpenStudy (uri):

-3x-9=-3x-3-5 -3x+3x=-3-5+9

OpenStudy (uri):

-3x+3x will be cancelled.

OpenStudy (anonymous):